OpenStudy (anonymous):

The perimeter of a rectangle is 52. If its length is 2 more than 3 times its width, find the dimensions of the rectangle.

OpenStudy (dan815):

hello so if x = length of rectangle and y=width of rectangle then 2x+2y=52, as there are 2 lines of each and 2 more than 3 times its width is length so x=3y+2 ---------------- solve these 2 equations 2x+2y=52 x=3y+2

OpenStudy (paki):

@dan815 nice explanation man...

OpenStudy (anonymous):

first one is 26

OpenStudy (dan815):

width or length?

OpenStudy (dan815):

if you found one you can plug in and find the other

OpenStudy (dan815):

Are you familiar with solving system of equations?

OpenStudy (anonymous):

width

OpenStudy (anonymous):

its 26 @dan815 ?

OpenStudy (dan815):

Solve this to find out: 2x+2y=52 x=3y+2 ------------

OpenStudy (dan815):

since you know x in terms of y, sub that into equation 1 and solve for y

OpenStudy (dan815):

2 (x) +2y=52 2(3y+2)+2y=52

OpenStudy (dan815):

2(3y+2)+2y=52 6y+4+2y=52 8y=48 y=6

OpenStudy (dan815):

x=3y+2 =3(6)+2